Circular No. 6448

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 P2 R. H. McNaught, Anglo-Australian Observatory, reports the discovery

by Kenneth S. Russell of a new comet on U.K. Schmidt plates taken by Fred G. Watson. The comet has a strong condensation (m_{2} ∼ 17) within a 20′′ coma, and a 20′ tail in p.a. 240^{o}. The following positions were measured by McNaught:

1996 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 10.74014  1^{h}40^{m}24.^{s}85  −29^{o}51′58′′.8
10.80264  1 40 25.78  −29 52 32.6
11.75125  1 40 39.80  −30 02 11.9  13
11.79292  1 40 40.36  −30 02 35.5

COMET C/1995 O1 (HALE-BOPP) R. W. Russell and D. Lynch, The Aerospace Corporation; M. Hanner,

Jet Propulsion Laboratory; and M. Sitko, University of Cincinnati, report observations of C/1995 O1 on July 22.3 UT with the Infrared Telescope Facility and the Aerospace infrared spectrograph between 3 and 13.5 μm. The 8- to 13.5-μm spectrum showed a strong silicate emission feature (70 percent above the continuum), similar to those seen in comets 1P/Halley, C/1987 P1 (Bradfield), and C/1990 K1 (Levy). After division by a blackbody at the continuum temperature, the feature was trapezoidal with a roughly-flat top from ∼ 9.9 to 11.5 μm and showed no significant additional spectral structure. The dust-continuum color temperature (from fitting a gray body to the 8- and 13-μm regions) was 200 ± 10 K, or roughly 50 K warmer than the expected radiative equilibrium temperature of 145 K for a blackbody at 3.7 AU from the sun. Due to the cold continuum shape, no detection of the comet was made in the 3- to 5-μm region. Brightnesses of the comet in the 3′′.4-diameter beam at 8.0, 10.2, and 13.0 μm were 2, 6, and 5 × 10^{−18} W cm^{−2} μm^{−1} (magnitudes 5.4, 3.2, and 2.4, respectively).

COMET C/1996 N1 (BREWINGTON) Total visual magnitude estimates (cf. IAUC 6435): July 27.27 UT, 8.8

(C. E. Spratt, Victoria, BC, 0.20-m reflector); Aug. 4.90, 8.6 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 7.24, 8.2 (Spratt); 9.24, 7.9 (Spratt); 10.93, 7.8 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector); 11.24, 7.8 (Spratt); 11.90, 7.6 (Bouma, 15×80 binoculars).

1996 August 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green